Abdul made $306 for 18 hours of work.
At the same rate, how much would he make for 13 hours of work?

Found 2 solutions by JulietG, stanbon:
Answer by JulietG(1812)

Abdul makes 306/18, or $17 an hour. He must be a skilled worker!
$17 * 13 hours = $221

Answer by stanbon(75887)

Abdul made $306 for 18 hours of work.
At the same rate, how much would he make for 13 hours of work?
---------
Use a Proportion:
x/13 = 306/18
----
x = 13(306/18) = $221.00
-------------------------
Cheers,
Stan H.
--------------